# SvANet: A Scale-variant Attention-based Network for Small Medical Object Segmentation

### v3

@@ Abstract
Early detection and accurate diagnosis can predict the risk of malignant disease transformation, thereby increasing the probability of effective treatment. A mild syndrome with small infected regions is an ominous warning and is foremost in the early diagnosis of diseases. Deep learning algorithms, such as convolutional neural networks (CNNs), have been used to segment natural or medical objects, showing promising results. However, analyzing medical objects of small areas in images remains a challenge due to information losses and compression defects caused by convolution and pooling operations in CNNs. These losses and defects become increasingly significant as the network deepens, particularly for small medical objects. To address these challenges, we propose a novel scale-variant attention-based network (SvANet) for accurate small-scale object segmentation in medical images. The SvANet consists of Monte Carlo attention, scale-variant attention, and vision transformer, which incorporates cross-scale features and alleviates compression artifacts for enhancing the discrimination of small medical objects. Quantitative experimental results demonstrate the superior performance of SvANet, achieving 96.12%, 96.11%, 89.79%, 84.15%, 80.25%, 73.05%, and 72.58% in mean Dice coefficient for segmenting kidney tumors, skin lesions, hepatic tumors, polyps, surgical excision cells, retinal vasculatures, and sperms, which occupy less than 1% of the image areas in KiTS23, ISIC 2018, ATLAS, PolypGen, TissueNet, FIVES, and SpermHealth datasets, respectively.
@@
Small object detection, medical image segmentation, attention mechanisms, Monte Carlo method, vision transformer
@@ I Introduction
It is essential to detect and diagnose diseases or conditions at their earliest stages, often prior to the manifestation of symptoms. Early detection can aid in the discovery of potential diseases, such as precancerous stages, leading to higher survival rates of patients [1, 2, 3, 4]. For instance, Sung et al. [5] found that the five-year survival rate for melanoma could reach 99% if diagnosed and treated early, compared to only 27% if detected in the late stage. In the early stages of diseases like glaucoma [1], skin cancer [2], colorectal cancer [3], hepatocellular carcinoma [6], renal cancer [4], etc., the affected areas are comparatively small and difficult to detect. The morphometrics of these affected areas is believed to reflect the risk (e.g., cancer precursors) and progression of diseases [1, 2, 3, 6, 4, 7]. Accurately delineating the boundaries of lesions is crucial for their complete resection. Cell-level imaging analysis is also a cutting-edge field with various clinical applications, such as tumor resection analysis [7] and in vitro fertilization [8]. However, examining cells can be challenging due to differences in size, morphology, and density, especially on a small scale.
Figure 1: Architecture of the scale-variant attention-based network (SvANet). Monte Carlo attention (MCAttn) and MCAttn-based bottleneck (MCBottleneck) modules, positioned in the top-left dashed boxes, learn global and local features, respectively. Cross-scale guidance and scale-variant attention (SvAttn) techniques, shown in the top-middle dashed boxes, integrate low-level and high-level feature maps. Convolution with vision transformer by assembling tensors (AssemFormer), located in top-right dashed boxes, synergistically correlates local and global features, combining the strengths of convolutional neural networks and vision transformers.
Various methodologies, including biomarkers, screening programs, medical imaging, genomics, and epidemiology, are commonly used for early disease detection and diagnosis. Recently, deep learning algorithms have demonstrated significant potential in enhancing diagnostic efficiency by reducing costs and interpreting images from various imaging modalities, such as ophthalmoscopy (Oph), dermatoscopy (Derm), colonoscopy (COL), magnetic resonance imaging (MRI), computerized tomography (CT), whole slide imaging (WSI), microscopy (Microsc), electron microscopy (EM), and X-ray [9, 10, 11, 12, 13, 14, 15, 16]. Compared to experienced medical specialists, deep learning algorithms have achieved better or comparable results in medical image segmentation. However, these methods often neglect the small-sized medical objects.
Medical objects that occupy less than 10% of the area pose a significant challenge in medical imaging due to their subtle texture and morphology. Detecting the malignant potential of polyp lesions smaller than 10 mm, for example, is a challenging task [3]. In reality, a large number of images from various modalities contain numerous lesions that occupy less than 10% of the total image area [1, 2, 3, 6, 4, 7], as detailed in Tab. I.
@@
Deep learning algorithms, which employ convolution and pooling, can result in the loss of details for small objects, leading to noticeable compression artifacts. Therefore, it is paramount to develop practical methods for detecting small medical objects. To address the issue of diminished image resolution and information loss, four main strategies have been developed. These strategies include upscaling input data [17], expanding network variants [9, 10, 11, 14, 12, 13, 15, 16], tuning loss functions [18, 19, 15], and post-processing [15]. Most recently, the vision transformer (ViT) has been introduced to process sequences of image patches to learn the inter-patch representations, which has shown noticeable potential in aggregating and preserving the features of small objects [16].
This study presents SvANet, a novel scale-variant attention-based architecture designed to segment medical objects. The primary goal of SvANet is to mitigate disease progression and enable timely interventions. The key contributions of this study are highlighted below:
@@
  • We propose SvANet, a new network that utilizes two novel attention mechanisms and a vision transformer to identify small medical objects. To the best of our knowledge, this is the first study to systematically analyze small medical objects across various medical image modalities (i.e., Oph, Derm, COL, MRI, CT, WSI, and Microsc) and diverse object types (i.e., retinal vessels, skin lesions, polyps, livers, kidneys, tumors, tissue cells, and sperms).
  • We introduce the Monte Carlo attention (MCAttn) module, which generates attention maps at different scales in a single stage by using agnostic pooling output sizes. MCAttn learns the object relations and spatial information of small medical objects with consideration of both their position and morphology.
@@
  • We develop the scale-variant attention (SvAttn) method, which captures the positional and morphological essence of small medical objects by generating attention maps based on the progressively compressed feature maps.
@@
  • To further enhance the recognition of small-scale medical objects, we present AssemFormer, a method that combines convolution with a vision transformer by assembling tensors. This approach enables the incorporation of both local spatial hierarchies and inter-patch representations, providing a comprehensive understanding of the image data.
@@ II-A Medical Object Segmentation
Surface structures, shapes, and sizes are critical in characterizing medical objects. The morphometrics are collected by various devices from different patients, making them complex and challenging to analyze. In recent years, deep learning techniques in computer vision have obtained promising performance for medical image segmentation. One widely adopted structure for analyzing medical images is the encoder-decoder-based construction, introduced by Long et al. in 2015 [20]. This approach involves extracting derived features from an encoder and using a decoder to generate the final segmentation mask. Building upon the encoder-decoder structure, Ronneberger et al. [9] introduced “U-shaped” architectures, which connect the limbs by using convolution (UNet) to disseminate information for segmenting tumor cells or general objects. To further enhance the fusion of multi-scale features in analyzing medical images across CT, MRI, and EM modalities, Zhou et al. [10] introduced UNet++, an extension of UNet that incorporates densely connected links.
@@ II-B Small Medical Object Segmentation
The convolution and pooling operations in deep learning algorithms compress input data, thus damaging the morphological characteristics of medical objects. To mitigate information loss when reducing image resolution, one common method is to upscale the input images for generating high-resolution feature maps of small objects [17]. However, this method can be time-consuming during training and testing due to the need for image augmentation and feature dimension enlargement. Another promising method is to expand network variants by incorporating techniques such as atrous convolution [22], skip connections [9, 10], feature pyramids [23, 14, 24], multicolumns [25, 14, 13, 15], or attention mechanisms [11, 14, 12, 13, 15, 16], which captures cross-scale features and contributes to magnify small objects. For example, Zhao et al. [23] introduced the pyramid scene parsing network (PSPNet), which employs pyramid pooling and concatenates up-sampled features from multiple scales to improve context feature learning. Lou et al. [12] proposed a context axial reverse attention network (CaraNet) to detect small polyps and brain tumors with less than 5% size ratios. However, CaraNet lacks sufficient interpretability regarding its practicality for segmenting small medical objects, appearing more as a general design suited to the segmentation task.
Designing new loss functions is another practical way to boost small object identification. Guo et al. [18] proposed a loss function that adopts the boundary pixel’s neighbors to enhance the small object segmentation. Besides, Pan et al. [15] combined six different loss functions for nuclei diagnosis. However, the disadvantage of replacing the loss function is that it may not be semantically understandable [18, 19] or it can increase the computational complexity [15]. Post-processing, such as the watershed algorithm [15], can also be employed to enhance small object segmentation. However, post-processing is a distinct step from the segmentation model, and the network cannot adjust its weights to the post-processing results.
@@ II-C Attention Mechanisms
The attention mechanism is extensively employed in semantic segmentation to prioritize salient features. Various approaches have been proposed to incorporate attention in different ways. Hu et al. [27] applied the squeeze-excitation (SE) method to generate channel attention for learning semantic representations. Zhou et al. [13] employed channel attention to capturing boundary-aware features for enhancing polyp segmentation. To further extract spatial information, Woo et al. [28] combined channel attention with spatial attention in the convolution block attention module (CBAM). Hou et al. [29] further advanced CBAM by introducing coordinate attention (CoorAttn), which utilizes channel-wise average pooling to generate attention maps. Reverse attention is another practical method to mine boundary cues. PraNet [11] extracted fine-grained details by removing the estimated polyps regions using boundary information. Lou et al. [12] enhanced PraNet by decomposing attention maps along height and width axes. Moreover, self-attention is an effective attention scheme to obtain dependencies and relationships within input data. He et al. [30] proposed a fully-transformer-based network that amalgamated spatial pyramid theory and vision transformer to identify skin lesions. However, the vanilla vision transformer lacks inherent bias and is susceptible to perturbations [31]. Zhang et al. [14] and Pan et al. [15] employed self-attention to improve the feature correlations in their CNN-based network for polyps and nuclei examination, respectively. To obtain long-range information when segmenting cell nuclei, Hörst et al. [32] replaced the CNN encoder with a transformer block in the UNet architecture in their CellViT model.
@@ III-A Overall Framework
This section introduces the scale-variant attention-based network (SvANet), specifically designed to segment small medical objects. The SvANet model, schematically depicted in Fig. 1, comprises four main components: Monte Carlo attention in Sec. III-B, cross-scale guidance in Sec. III-C, scale-variant attention in Sec. III-D, and the convolution with vision transformer in Sec. III-E.
Human eyes examine objects by altering the shape of crystalline lenses, thereby creating a detailed visual construction of the object’s characteristics. When viewing an object at various distances, changes in the lens enhance the capture of object details (as per SvAttn and AssemFormer) and changes in location behaviors (as per cross-scale guidance and SvAttn). Drawing inspiration from this phenomenon, we present SvANet by integrating MCAttn, cross-scale guidance, SvAttn, and AssemFormer to mitigate compression defects and enhance the analysis of small medical objects.
@@
To improve the identification of small medical objects and maintain the plug-and-play design of the encoder division, a Monte Carlo attention (MCAttn) based bottleneck (MCBottleneck) is integrated into each encoder stage, depicted by the blue arrows and the blue dashed block in Fig. 1. Preserving the features of tiny medical objects, such as sperms and retinal vessels, becomes challenging after multiple pooling or strided convolution operations. In this study, cross-scale feature maps are applied to guide the latter stages in learning features of small medical objects, represented by the orange arrows in Fig. 1. To improve the identification of object features and their correlations, the correlated features are further processed by convolutions integrated with transformers (i.e., green blocks in Fig. 1). To better comprehend the role of MCAttn, cross-scale guidance, and SvAttn in small medical object segmentation, we examined the feature maps as presented by Fig. 2 and Fig. 3. Due to page limitations, we selected FIVES, ISIC 2018, KiTS23, and SpermHealth as example datasets in Fig. 2 and Fig. 3. To simplicity, we chose to present the outputs from MCBottleneck in stage four and two cross-scale guidance in Fig. 2 and Fig. 3.
@@
Additionally, in Fig. 1, each Conv3×3↓2, represented by black blocks, contains a single 3×3 convolution with a stride of 2 (strided convolution). Every TConv3×3↓2, denoted by grey blocks, consists of three convolution units: a 1×1 convolution, a 3×3 transposed strided convolution, and a 1×1 convolution. Besides, each MCBottleneck consists of a 3×3 convolution, a 1×1 convolution, and an MCAttn module (see Sec. III-B), followed by a 1×1 convolution and an AssemFormer block (see Sec. III-E). The MCBottleneck serves as a compression point in the network, narrowing the tensor channels before expanding them to extract salient features by compressing the input information, resembling a “bottleneck” in information theory. To expand the receptive field and capture features at multiple scales, atrous spatial pyramid pooling (ASPP) [22] is integrated after the final stage of our model.
III-B Monte Carlo Attention
Inspired by this phenomenon, the Monte Carlo attention (MCAttn) module uses a random-sampling-based pooling operation to generate scale-agnostic attention maps. This enables the network to capture relevant information across different scales, enhancing its ability to identify small medical objects.
@@ III-B Monte Carlo Attention
Figure 2: Output feature maps from MCBottleneck without or with an attention mechanism. (a) Input image, (b) ground truth, and feature map generated (c) without using attention mechanism and with using (d) SE, (e) CBAM, (f) CoordAttn, and (g) Monte Carlo attention (MCAttn). Example images in odd and even rows include ultra-small and small medical objects, respectively.
In Fig. 1, the MCAttn, as presented by the purple block, generates attention maps by randomly selecting a 1×1 attention map from three scales: 3×3, 2×2, and 1×1 (pooled tensors). In conventional methods such as squeeze-excitation (SE), global-average pooling is used to acquire a 1×1 output tensor, which helps calibrate the inter-dependencies between channels [27]. However, this approach has limited capacity to exploit cross-scale correlations. To address this limitation, MCAttn calculates the attention maps from features across three scales, thereby enhancing long-range semantic interdependences. Given an input tensor, $x$, the output attention map of MCAttn, denoted as $\mathcal{A}_{\text{m}}(x)$, is computed as follows:
$\mathcal{A}_{\text{m}}(x)=\sum_{i=1}^{n}P_{1}(x,i)f(x,i)$ (1)
The Monte Carlo sampling method described in Eq. 1 allows for the random selection of association probabilities, enabling the extraction of both local information (e.g., angle, edge, and color) and context information (e.g., whole image texture, spatial correlation, and color distribution). In Fig. 2c, the second to the fourth rows and the final row illustrate that MCBottleneck, without using a attention mechanism, struggles to detect the retinas and nevi and often overlooks several sperms. Conversely, when attention mechanisms like SE, CBAM, and CoordAttn are used, localization of densely occupied regions (e.g., optic disc, kidneys, and sperms) is enhanced compared to when no attention mechanism is used. However, sparse regions, such as retinal vasculatures and nevus centres, are often underlooked, especially the ultral-small ones, as shown in Fig. 2c-f. Instead, using MCAttn in MCBottleneck, as depicted in Fig. 2cg, enhances the discernibility of the morphology and precise location of both ultra-small and small medical objects compared to when MCAttn is not used. For instance, in Fig. 2, moving downwards, MCBottleneck coupled with MCAttn emphasizes more apparent retinal vessels for glaucoma, sharper boundaries of nevi, and more perceptible morphology of kidneys, cysts, and sperms. MCAttn also accentuates other medical objects of interest, such as retinas, nevi, kidneys, and sperms, as shown in Fig. 2g.
III-C Cross-scale Feature Guidance

### v4

# Exploiting Scale-Variant Attention for Segmenting Small Medical Objects

@@ Abstract
Early detection and accurate diagnosis can predict the risk of malignant disease transformation, thereby increasing the probability of effective treatment. Identifying mild syndrome with small pathological regions serves as an ominous warning and is fundamental in the early diagnosis of diseases. While deep learning algorithms, particularly convolutional neural networks (CNNs), have shown promise in segmenting medical objects, analyzing small areas in medical images remains challenging. This difficulty arises due to information losses and compression defects from convolution and pooling operations in CNNs, which become more pronounced as the network deepens, especially for small medical objects. To address these challenges, we propose a novel scale-variant attention-based network (SvANet) for accurately segmenting small-scale objects in medical images. The SvANet consists of scale-variant attention, cross-scale guidance, Monte Carlo attention, and vision transformer, which incorporates cross-scale features and alleviates compression artifacts for enhancing the discrimination of small medical objects. Quantitative experimental results demonstrate the superior performance of SvANet, achieving 96.12%, 96.11%, 89.79%, 84.15%, 80.25%, 73.05%, and 72.58% in mean Dice coefficient for segmenting kidney tumors, skin lesions, hepatic tumors, polyps, surgical excision cells, retinal vasculatures, and sperms, which occupy less than 1% of the image areas in KiTS23, ISIC 2018, ATLAS, PolypGen, TissueNet, FIVES, and SpermHealth datasets, respectively.
@@
Small object detection, medical image segmentation, attention mechanisms, Monte Carlo method, vision transformer.
@@ I Introduction
It is essential to detect and diagnose diseases or conditions at their earliest stages, often prior to the manifestation of symptoms. Early detection can aid in the discovery of potential diseases, such as precancerous stages, leading to higher survival rates of patients [1, 2, 3, 4]. For instance, Sung et al. [5] found that the five-year survival rate for melanoma could reach 99% if diagnosed and treated early, compared to only 27% if detected in the late stage. In the early stages of diseases like glaucoma [1], skin cancer [2], colorectal cancer [3], hepatocellular carcinoma [6], renal cancer [4], etc., the pathological areas are comparatively small and difficult to detect. The morphometrics of these infected areas is believed to reflect the risk (e.g., cancer precursors) and progression of diseases [1, 2, 3, 6, 4, 7]. Accurately delineating the boundaries of lesions is crucial for their complete resection. Cell-level imaging analysis is also a cutting-edge field with various clinical applications, such as tumor resection analysis [7] and in vitro fertilization [8]. However, examining cells can be challenging due to differences in size, morphology, and density, especially on a small scale.
@@
Medical objects that occupy less than 10% of the area pose a significant challenge in medical imaging due to their subtle texture and morphology. Detecting the malignant potential of polyp lesions smaller than 10 mm, for example, is a challenging task [3]. In reality, a large number of images from various modalities contain numerous lesions that occupy less than 10% of the total image area [1, 2, 3, 6, 4, 7], as detailed in Tab. I. Deep learning algorithms, which employ convolution and pooling, can result in the loss of details for small objects, leading to noticeable compression artifacts. Therefore, it is paramount to develop practical methods for detecting small medical objects. To address the issue of diminished image resolution and information loss, four main strategies have been developed. These strategies include upscaling input data [9], expanding network variants [10, 11, 12, 13, 14, 15, 16, 17, 18], tuning loss functions [19, 20, 17], and post-processing [17].
Figure 1: Illustration of the intuitions of the core components of the proposed methods. The example image depicts a skin lesion.
Attention is a fundamental process for conscious perception and effective interaction with the environment. In deep learning, the attention mechanism is an efficient method to extend network variants [12, 15, 16, 17, 18]. However, small medical objects pose unique challenges: they not only lack sufficient pixels and information for straightforward local representation extraction, but their relatively small size (e.g., occupying less than 1% of the images) makes them difficult to capture using global operations such as global average pooling and multi-head self-attention. Animal eyes provide a useful analogy for addressing these challenges by altering the shape of their crystalline lenses, thereby creating a detailed visual construction of objects and capturing changes in location and size at various distances. Inspired by these biological mechanisms, we introduce a scale-variant attention (SvAttn) method within a cross-scale guidance module for “tracing” the behavior of small medical objects by using cross-level features, as demonstrated in Fig. 1a.
Figure 2: Architecture of the scale-variant attention-based network (SvANet). Cross-scale guidance and scale-variant attention (SvAttn) techniques, depicted in the top-left dashed boxes, integrate low-level and high-level feature maps to trace the alterations in the shape and location of small medical objects. The modules Monte Carlo attention (MCAttn) and MCAttn-based bottleneck (MCBottleneck), positioned in the top-middle dashed boxes, along with assembly-based convolutional vision transformer (AssemFormer) in the top-right dashed boxes, synergistically correlate local and global features to capture intricate object details.
Furthermore, traditional attention mechanisms in deep learning typically produce a fixed-dimension attention map [16, 21, 22, 23], which is often inadequate to analyze medical images. This limitation arises because these methods focus primarily on central features, neglecting the rich contextual information in the background, which is crucial for clinical interpretation. For instance, in an abdominal slice image, the standard positional relationships among various organs (e.g., stomach, liver, kidneys, spleen, and bone marrow) aid in accurately locating objects of interest within a narrower range. Drawing inspiration from this observation, we introduce the Monte Carlo attention (MCAttn) and an assembly-based convolutional vision transformer (AssemFormer) to address compression defects for “detailing” features of small medical objects from different scales, as illustrated by Fig. 1b.
The key contributions of this study are highlighted below:
@@
  • We propose SvANet, a new network that utilizes two novel attention mechanisms and a vision transformer to identify small medical objects. To the best of our knowledge, this is the first study to systematically analyze small medical objects across seven medical image modalities and diverse object types (i.e., retinal vessels, skin lesions, polyps, livers, kidneys, tumors, tissue cells, and sperms).
@@
  • We introduce the SvAttn method, which captures the positional and morphological essence of small medical objects by generating attention maps based on the progressively compressed feature maps.
  • We develop the MCAttn module, which generates attention maps at different scales in a single stage by using agnostic pooling output sizes. MCAttn learns the object relations and spatial information of small medical objects with consideration of both their position and morphology.
@@
  • We present AssemFormer, which enables the incorporation of both local spatial hierarchies and inter-patch representations, providing a comprehensive understanding of the image data.
@@ II-A Medical Object Segmentation
Surface structures, shapes, and sizes are critical in characterizing medical objects. The morphometric data, collected from various devices and different patients, present a complex and challenging landscape for analysis. In recent years, deep learning algorithms have shown remarkable potential in enhancing diagnostic accuracy, reducing costs, and interpreting images of diverse medical objects across various imaging modalities. Thes modalities include ophthalmoscopy (Oph), dermatoscopy (Derm), colonoscopy (COL), magnetic resonance imaging (MRI), computerized tomography (CT), whole slide imaging (WSI), microscopy (Microsc), electron microscopy (EM), X-ray, and etc. [10, 11, 12, 13, 14, 15, 16, 17, 18].
One widely adopted structure for analyzing medical images is the encoder-decoder-based construction, introduced by Long et al. [24] in 2015. This approach involves extracting derived features from an encoder and using a decoder to generate the final segmentation mask. Building upon the encoder-decoder structure, Ronneberger et al. [10] introduced “U-shaped” architectures, which connect the limbs by using convolution (UNet) to disseminate information for segmenting tumor cells or general objects. To further enhance the fusion of multi-scale features in analyzing medical images across CT, MRI, and EM modalities, Zhou et al. [11] introduced UNet++, an extension of UNet that incorporates densely connected links. In addition, Isensee et al. [13] broadened the application of UNet from 2D to 3D medical imaging by self-adaptive configurations (nnUNet).
@@ II-B Small Medical Object Segmentation
The convolution and pooling operations in deep learning algorithms compress input data, thus damaging the morphological characteristics of medical objects. To mitigate information loss when reducing image resolution, one common method is to upscale the input images for generating high-resolution feature maps of small objects [9]. However, this method can be time-consuming during training and testing due to the need for image augmentation and feature dimension enlargement. Another promising method is to expand network variants by incorporating techniques such as atrous convolution [26], skip connections [10, 11, 13], feature pyramids [27, 28], multicolumns [29, 14, 16, 17], or attention mechanisms [12, 15, 16, 17, 18], which captures cross-scale features and contributes to magnify small objects. For example, Zhao et al. [27] introduced the pyramid scene parsing network (PSPNet), which employs pyramid pooling and concatenates up-sampled features from multiple scales to improve context feature learning. Lou et al. [15] proposed a context axial reverse attention network (CaraNet) to detect small polyps and brain tumors with less than 5% size ratios. However, CaraNet lacks sufficient interpretability regarding its practicality for segmenting small medical objects, appearing more as a general design suited to the segmentation task.
Designing new loss functions is another practical way to boost small object identification. Guo et al. [19] proposed a loss function that adopts the boundary pixel’s neighbors to enhance the small object segmentation. Besides, Pan et al. [17] combined six different loss functions for nuclei diagnosis. However, the disadvantage of replacing the loss function is that it may not be semantically understandable [19, 20] or it can increase the computational complexity [17]. Post-processing, such as the watershed algorithm [17], can also enhance small object segmentation. However, post-processing is a distinct step from the segmentation model, and the network cannot adjust its weights to the post-processing results.
@@ II-C Attention Mechanisms
The attention mechanism is extensively employed in semantic segmentation to prioritize salient features. Various approaches have been proposed to incorporate attention in different ways. Hu et al. [21] applied the squeeze-excitation (SE) method to generate channel attention for learning semantic representations. Zhou et al. [16] employed channel attention to capturing boundary-aware features for enhancing polyp segmentation. To further extract spatial information, Woo et al. [22] combined channel attention with spatial attention in the convolution block attention module (CBAM). Hou et al. [23] further advanced CBAM by introducing coordinate attention (CoorAttn), which utilizes channel-wise average pooling to generate attention maps. Reverse attention is another practical method to mine boundary cues. PraNet [12] extracted fine-grained details by removing the estimated polyps regions using boundary information. Lou et al. [15] enhanced PraNet by decomposing attention maps along height and width axes. Relatively small feature maps (ranging from 1×1 to 44×44) were employed to bridge area and boundary cues, which may not adequately capture the structural details of minuscule objects [21, 12, 18].
Moreover, self-attention is an effective attention scheme to obtain dependencies and relationships within input data. Based on self-attention mechanisms, the vision transformer (ViT) has been introduced to process sequences of image patches to learn the inter-patch representations, which has shown noticeable potential in aggregating and preserving the features of small objects [18]. He et al. [31] proposed a fully-transformer-based network that amalgamated spatial pyramid theory and vision transformer to identify skin lesions. However, the vanilla vision transformer lacks inherent bias and is susceptible to perturbations [32]. Zhang et al. [14] and Pan et al. [17] employed self-attention to improve the feature correlations in their CNN-based network for polyps and nuclei examination, respectively. To obtain long-range information when segmenting cell nuclei, Hörst et al. [33] replaced the CNN encoder with a transformer block in the UNet architecture in their CellViT model.
@@ III-A Overall Framework
This section introduces the scale-variant attention-based network (SvANet), specifically designed to segment small medical objects. The SvANet model, schematically depicted in Fig. 2, comprises four main components: cross-scale guidance in Sec. III-B, scale-variant attention in Sec. III-C, Monte Carlo attention in Sec. III-D, and the convolution with vision transformer in Sec. III-E.
@@
Preserving the features of tiny medical objects, such as sperms and retinal vessels, becomes challenging after multiple pooling or strided convolution operations. For example, after two strided convolutions, a sperm may be reduced to being represented by only one or two pixels in the image. In this study, cross-scale feature maps are applied to guide the latter stages in learning the features of small medical objects, as indicated by the orange arrows in Fig. 2. The SvAttn and cross-scale guidance are primarily designed to track feature changes, particularly downsizing. Meanwhile, MCAttn and AssemFormer distill multi-scale attention maps for improved contextual feature learning. To better comprehend the roles of cross-scale guidance, SvAttn, MCAttn, and AssemFormer in small medical object segmentation, we examined the feature maps as shown in Fig. 4, Fig. 5, and Fig. 3. Due to page limitations, we selected FIVES, ISIC 2018, KiTS23, and SpermHealth datasets to visualize feature maps. To simplicity, we chose to present the outputs from the MCBottleneck in stage four (Fig. 4) and two cross-scale guidance correlations (Fig. 3).
@@
Additionally, in Fig. 2, each Conv3×3↓2, represented by black blocks, contains a single 3×3 convolution with a stride of 2 (strided convolution). Every TConv3×3↓2, denoted by grey blocks, consists of three convolution units: a 1×1 convolution, a 3×3 transposed strided convolution, and a 1×1 convolution. The MCBottleneck serves as a compression point in the network, narrowing the tensor channels before expanding them to extract salient features by compressing the input information, resembling a “bottleneck” in information theory. To expand the receptive field and capture features at multiple scales, atrous spatial pyramid pooling (ASPP) [26] is integrated after the final stage of our model.
@@
Figure 3: Output feature maps from cross-scale guidance without or with scale-variant attention (SvAttn). These feature maps are generated from $g(x_{s},t)$, which integrates features at different scales. Example images in odd and even rows include ultra-small and small medical objects, respectively (GT: ground truth).
III-B Cross-Scale Feature Guidance